An object has kinetic energy of 18j. If it's speed is 3m/s,what is its mass
Answer #1

formula for kinemtic energy is

KE = 1/2 m v^2

18 = 1/2 * m ( 3)^2

m = 2 ( 18)/ 3^2

m = 4 kg